Circular No. 7501

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 S3 On Oct. 1, B. W. Koehn communicated his measurements of a comet,

later reported as discovered by B. A. Skiff on images taken on Sept. 29 by M. E. Van Ness in the course of the LONEOS program at Lowell Observatory. Skiff described a nearly circular coma of diameter 15′′ with moderate condensation. Following tentative linkage by B. G. Marsden, Center for Astrophysics, to one (mag ∼ 19) of ∼ 2000 asteroidal objects recorded on Sept. 20 by LINEAR, a tentative ephemeris was provided on the NEO Confirmation Page. This linkage was confirmed by observations obtained on Oct. 2 by J. G. Ries with the 0.76-m reflector at McDonald Observatory. Skiff adds that observations by L. H. Wasserman with the 1.1-m telescope at Lowell Observatory on Oct. 2 showed a 14′′ × 11′′ coma, elongated eastwest. Positions, orbit, and ephemeris follow:

2000 UT α_{2000} δ_{2000} m_{1} Observer
Sept.20.33024 0^{h}59^{m}03.^{s}38 −21^{o}43′42′′.8  LINEAR
20.34573 0 59 02.57 −21 43 39.5  _{”}
20.36106 0 59 01.91 −21 43 34.9  _{”}
20.37617 0 59 01.20 −21 43 30.3  _{”}
20.39095 0 59 00.39 −21 43 26.3  _{”}
29.26958 0 51 46.10 −21 01 00.7 17.7 Van Ness
29.27891 0 51 45.60 −21 00 57.2  _{”}
29.28826 0 51 45.18 −21 00 53.8  _{”}
29.29753 0 51 44.66 −21 00 51.4  _{”}
Oct. 2.26013 0 49 16.18 −20 42 59.5  Ries
2.26713 0 49 15.84 −20 42 57.1  _{”}
2.27394 0 49 15.46 −20 42 54.7  _{”}
T = 2000 July 15.867 TT  ω = 298.^{o}162
e = 0.77299  Ω = 41.137 }2000.0
q = 2.66273 AU  i = 25.179
a = 11.72976 AU n^{o} = 0.024534 P = 40.17 years
2000TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 23 0^{h}56.^{m}93 −21^{o}32.^{′}6 1.799 2.738 154.^{o}4 9.^{o}1 17.6
Oct. 3 0 48.65 −20 38.2 1.821 2.761 155.2  8.7 17.7
13 0 40.62 −19 22.8 1.868 2.787 151.6  9.8 17.8
23 0 33.71 −17 48.3 1.941 2.815 145.1 11.7 17.9
2000 October 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green